OpenStudy (anonymous):

evaluate the iterated integral: (from 0 to ln3),(from 1 to ln2) e^(2x+4y)dydx

OpenStudy (anonymous):

\[\int\limits_{0}^{\ln3}\int\limits_{1}^{\ln2} e^{(2x+4y)}dydx\]

OpenStudy (anonymous):

\[\int\limits_{0}^{\ln3}\ \left( \int\limits_{1}^{\ln4}e^{(2x+4y)}dy \right)dx\]

OpenStudy (anonymous):

inside the parentheses would be this? \[\frac{ 1 }{ 4 }e^{2x+4(\ln2)}-\frac{ 1 }{ 4 }e^{2x+4}\]

OpenStudy (anonymous):

I entered the wrong upper limit in the parentheses - it is supposed to be ln2

ganeshie8 (ganeshie8):

Looks good

ganeshie8 (ganeshie8):

You could also split the double integral into a product of two single integrals

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{\ln3}\int\limits_{1}^{\ln2} e^{(2x+4y)}dydx \] \[\int\limits_{0}^{\ln3}e^{2x}\int\limits_{1}^{\ln2} e^{4y}dydx \] \[\left(\int\limits_{0}^{\ln3}e^{2x}dx \right)\cdot \left(\int\limits_{1}^{\ln2} e^{4y}dy\right) \]

ganeshie8 (ganeshie8):

that works whenever you can write the integrand as a product of form : "f(x)*g(y)"

OpenStudy (anonymous):

That looks much easier to deal with

OpenStudy (anonymous):

I messed up somewhere: \[\left( \frac{ 1 }{ 2 }e^{2(\ln3)}-\frac{ 1 }{ 2 }e^{2(0)} \right) * \left( \frac{ 1 }{ 4 }e^{4(\ln2)}-\frac{ 1 }{ 4 }e^{4(1)} \right)\ \]

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ 2 }\left( 9-1 \right) \right)*\left( \frac{ 1 }{ 4 } \left( 16-e \right)\right)\] \[\left( 4 \right)*\left( 4-\frac{ e }{ 4 } \right)\] \[16-e\]

OpenStudy (perl):

do you have any more questions on that

OpenStudy (anonymous):

It's not the right answer. I made a mistake and can't spot it.

zepdrix (zepdrix):

Oh oh oh I see.`

zepdrix (zepdrix):

\[\Large\rm e^{4(1)}\]For the lower limit on your y's

zepdrix (zepdrix):

\[\Large\rm \ne e^1\]

OpenStudy (anonymous):

oh! so that makes the end result 16-e^4?

zepdrix (zepdrix):

looks good!

OpenStudy (anonymous):

Thanks for catching that!

OpenStudy (jhannybean):

Good eye